# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ A.3.3 Optimized Information After M Has Learned from Training
$S_{M}(M,Tu_{m})=\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\}$ (A.36)
is a one-to-one enabling mapping from $S_{ou}(ou,Tu_{h})$ onto M, representing the state of M after completing its training.
A.3.4 Input Information for M
@@
$I_{q}=(oq,Tq_{h},S_{oq},Mq,Tq_{m},S_{Mq})$ (A.37)
Here, $oq$ denotes the object described by the input information, and $Tq_{h}$ is the time at which the object’s state occurs. The state $S_{oq}(oq,Tq_{h})$ can take two possible forms:
@@
    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1}\}$ (A.38)
    $\varphi q_{1}=\forall XBelongsTo^{2}(X,\mathbb{R}^{2})$ (A.39)
@@
    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1},\varphi q_{2}\}$ (A.40)
    $\varphi q_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.41)
    $\varphi q_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.42)
The carrier $Mq$ for $I_{q}$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$. The state $S_{Mq}(Mq,Tq_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,Tq_{h})$ onto $Mq$.
A.3.5 Information After M Processes Input, Produces Output, and Undergoes Further Optimization
There are two scenarios for M after it processes input and produces output: one corresponds to the first case in A.3.4, where only output is produced without further optimization of the model itself; the other corresponds to the second case in A.3.4, where, in the context of online learning, the model both produces output and undergoes further self-optimization. Since the latter is more comprehensive, we will focus on this scenario below.
In this case, the output information $Ir=(or,Tr_{h},S_{or},Mr,Tr_{m},S_{Mr})$ has an ontological state defined as:
$S_{or}(or,Tr_{h})$ $=\{\varphi r_{1}\}=\{\varphi o_{7}\land\varphi o_{8}\land\text{OutputAt}^{2}(Y,Tr_{h})\}$ (A.43)
The carrier state $S_{Mr}(Mr,Tr_{m})$ is a one-to-one enabling mapping from $S_{or}(or,Tr_{h})$ to the output device of M.
@@ A.3.5 Information After M Processes Input, Produces Output, and Undergoes Further Optimization
$S_{ov}(ov,Tv_{h})=\{\varphi v_{1},\varphi v_{6},\varphi v_{7},\varphi v_{8},\varphi v_{9},\varphi v_{10},\varphi v_{11}\}$ (A.45)
Here, $\varphi v_{1}$, $\varphi v_{6}$, and $\varphi v_{9}$ are identical to $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{9}$ respectively. For the remaining formulas, it suffices to systematically replace the parameters $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$ with the corresponding parameters $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$, $b_{i}^{(v)k}$, $W_{1}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(w)}$ and $b_{i}^{(w)k}$ to obtain the updated formulas, where $W_{k}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(2)}$ and $b_{i}^{(w)k}$ are the weights and biases of M after further optimization.
@@
$S_{M}(M,Tv_{m})=\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},\psi v_{10},\psi v_{11}\}$ (A.46)
is a one-to-one enabling mapping from $S_{ov}(ov,Tv_{h})$ onto M.
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Figure A1. Although M is a simple model, it contains all the essential components of a neural network, and the discussion has not relied on any special treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.

### v10

@@ C.3 Formal Representations of Information at Each Stage of Model M
$S_{M}\left(M,{Tu}_{m}\right)=\left\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\right\}$ (C.36)
is a one-to-one enabling mapping from $S_{ou}\left(ou,{Tu}_{h}\right)$ onto M, representing the state of M after completing its training.
(4) Input Information for M
@@
$Iq=\left\langle oq,{Tq}_{h},S_{oq},Mq,{Tq}_{m},S_{Mq}\right\rangle$ (C.37)
Here, $oq$ denotes the object described by the input information, and ${Tq}_{h}$ is the time at which the object's state occurs. The state $S_{oq}(oq,{Tq}_{h})$ can take two possible forms:
@@
$S_{oq}(oq,{Tq}_{h})=\left\{\varphi q_{1}\right\}$ (C.38)
$\varphi q_{1}=\forall X{BelongsTo}^{2}(X,\mathbb{R}^{2})$ (C.39)
@@
$S_{oq}(oq,{Tq}_{h})=\left\{\varphi q_{1},\varphi q_{2}\right\}$ (C.40)
$\varphi q_{1}=\forall X\forall Y_{t}({BelongsTo}^{2}(X,\mathbb{R}^{2})\land{BelongsTo}^{2}(Y_{t},\mathbb{R}^{2}))$ (C.41)
$\varphi q_{2}=\forall X\forall Y_{t}{Eq}^{2}(Target(X),Y_{t})$ (C.42)
The carrier $Mq$ for $Iq$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$. The state $S_{Mq}(Mq,{Tq}_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,{Tq}_{h})$ onto $Mq$.
(5) Information After M Processes Input, Produces Output, and Undergoes Further Optimization
There are two scenarios for M after it processes input and produces output: one corresponds to the first case in (4), where only output is produced without further optimization of the model itself; the other corresponds to the second case in (4), where, in the context of online learning, the model both produces output and undergoes further self-optimization. Since the latter is more comprehensive, we will focus on this scenario below.
In this case, the output information $Ir=\left\langle or,{Tr}_{h},S_{or},Mr,{Tr}_{m},S_{Mr}\right\rangle$ has an ontological state defined as:
$S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r_{1}\right\}=\left\{{\varphi o_{7}\land\varphi o_{8}\land OutputAt}^{2}(Y,{Tr}_{h})\right\}$ (C.43)
The carrier state $S_{Mr}(Mr,{Tr}_{m})$ is a one-to-one enabling mapping from $S_{or}\left(or,{Tr}_{h}\right)$ to the output device of M.
@@
$S_{ov}\left(ov,{Tv}_{h}\right)=\left\{\varphi v_{1},\varphi v_{6},\varphi v_{7},\varphi v_{8},\varphi v_{9},\varphi v_{10},\varphi v_{11}\right\}$ (C.45)
Here, $\varphi v_{1}$, $\varphi v_{6}$, and $\varphi v_{9}$ are identical to $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{9}$ respectively. For the remaining formulas, it suffices to systematically replace the parameters
$W_{k}^{(u)},w_{ij}^{(u)k},b_{k}^{(u)},b_{i}^{(u)k},W_{k}^{(v)},w_{ij}^{(v)k},b_{k}^{(v)},b_{i}^{(v)k}$
in $S_{ou}(ou,{Tu}_{h})$ with the corresponding parameters
$W_{k}^{(v)},w_{ij}^{(v)k},b_{k}^{(v)},b_{i}^{(v)k},W_{k}^{(w)},w_{ij}^{(w)k},b_{k}^{(w)},b_{i}^{(w)k}$
to obtain the updated formulas, where $W_{k}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(w)}$, and $b_{i}^{(w)k}$ are the weights and biases of M after further optimization.
@@
$S_{M}\left(M,{Tv}_{m}\right)=\left\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},{\psi v}_{10},\psi v_{11}\right\}$ (C.46)
is a one-to-one enabling mapping from $S_{ov}\left(ov,{Tv}_{h}\right)$ onto M.
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig. C.1. Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.